Assume a demand equation for goodx: where pown price of the good Q-quantity demanded Py price of a related good $3 Pz price of a different related good $200 Y = consumer income = $4,000/mo The quantity demanded as a function of the price can be written:
Answer #1

Replacing, py = 3, pz = 200 and Y = 4000 in the given equation, we get,

Q = 9 - 0.1 p - 3 + 0.01 (200) + 0.0005 (4000)

Q = 6 - 0.1p + 2 + 2 = 10 - 0.1 p

The quantity demanded as a function of the price can be written:

Q = 10 - 0.1p